OpenStudy (igster1112):

Need math help with problem like 4-2(-2 exponent)PLEASE HELP!!

OpenStudy (adamaero):

negative exponent is the same as doing 1/number eg: x^-2 = 1/(x^2)

OpenStudy (igster1112):

So what the answer?

OpenStudy (igster1112):

I need help on doing problems like this step by step.

OpenStudy (adamaero):

so you want help adding or subtracting fractions from whole numbers?

OpenStudy (adamaero):

\[5^{-2} = \frac{ 1 }{ 5^2 } = \frac{ 1 }{ 25 }\] \[10^{-1} = \frac{ 1 }{ 10 }\]

OpenStudy (igster1112):

Yes

OpenStudy (cj49):

4-1/2^2 4-1/4=?

OpenStudy (adamaero):

these are fundamental concepts you need to learn by doing yourself there are no handouts in life, read the textbook